OpenStudy (anonymous):

hey, I got the following problem. Why is cos(2x)=cos(x+x) ? I don't why. Please help me :)

OpenStudy (jdoe0001):

let's see x + x = 2x so... ahemm

OpenStudy (anonymous):

2x=x+x

OpenStudy (anonymous):

Oh my goodness, true, I did to much math today. thank you sou much guys!! :D

OpenStudy (anonymous):

Because x + x = 2x. Remember that cos is a function operating on whatever the argument is. So if we have two things that are equal, when we plug them in to the function, we get the same output.

OpenStudy (anonymous):

I got it, thanks!!!

OpenStudy (anonymous):

Do math every day.